the ine. The area of c labor dollars, while the audit f analysis and 1,700 Number of minutes 60,000 Facility rental 43,000 Per expense reports Direct
Overhead Rates Employee training per DL dollar per report/form Computing Minutes Emplovees per employee
per DL dollar
Cest Drivers head Ra training Computing Direct
Answer #1

Total overhead cost assigned :

Audit = 1,179,000 * 40% = 471,600

Tax= 626,000 *40% = 250,400

Activity Cost Pools Estimated overhead Expected use of cost drivers per activity Activity -based overhead rate
Employee training 252,700 1,805,000 Direct Labor dollars 0.14 per Direct Labor hour
Typing and secretarial 76,500 2,500 Reports/forms 30.6 per report/form
Computing 168,500 60,000 Minutes 2.8083 per minute
Facility Rental 143,000 40 Employees 3575 per employee
Travel 81,300 Direct Direct
Audit Tax
Activity Cost Pools Expected use of cost drivers Activity - based overhead rates Cost assigned Expected use of cost drivers Activity - based overhead rates Cost assigned
Employee training 1,179,000 0.14                          165,060 626,000 0.14               87,640
Typing and secretarial 800 30.6                             24,480 1,700 30.6               52,020
Computing 27,000 2.8083                             75,825 33,000 2.8083               92,675
Facility Rental 22 3575                             78,650 18 3575               64,350
Travel -                             56,000               25,300
Overhead costs assigned                          400,015             321,985